OpenStudy (anonymous):

If X is a uniformly distributed continuous random variable with possible values in the interval [-3,5] A)compute μx B)compute σ^2X C) Find P (|X| ≥ 1.5)

OpenStudy (anonymous):

Since \(X\) is uniformly distributed on \([-3,5]\), its probability density function is \[f(x)=\begin{cases} \dfrac{1}{5-(-3)}=\dfrac{1}{8}&\text{for }-3\le x\le5\\ 0&\text{otherwise} \end{cases}\] The mean \(\mu\) is \[E(X)=\int_{-\infty}^\infty f(x)~dx\] and the variance \(\sigma^2\) is \[\begin{align*} V(X)&=E(X^2)-[E(X)]^2\\ &=\int_{-\infty}^\infty x^2~f(x)~dx-\mu^2 \end{align*}\] For part (c), you're equivalently finding \[P(|X|\ge1.5)=P(-1.5\ge X\text{ or }X\le1.5)=P(X\ge1.5)+P(X\le-1.5)\] i.e. the total area of the shaded regions